OpenStudy (anonymous):

Multiply. (m^2 -2mn) (3mn + n^2)

sam (.sam.):

3m^3n-5m^2n^2-2mn^3

sam (.sam.):

\[3m ^{3}n-5m ^{2}n ^{2}-2mn ^{3}\]

OpenStudy (anonymous):

(m^2 -2mn) (n^2 + 3mn) = m^2n^2 + 3m^3n - 2mn^3 -6m^2n^2 = - 5m^2n^2 + 3m^3n - 2mn^3